OpenStudy (anonymous):

Which product is negative? They're all positive and I keep on re doing it again A. –3 • (–2) • (–4) • (–7) B. –2 • (–7) • 12 • (4) C. 4 • (–9) • (–3) • (–1) D. –6 • (–7) • (–8) • 0

OpenStudy (anonymous):

@whpalmer4

OpenStudy (anonymous):

I'll give a medal

OpenStudy (onepieceftw):

When you multiply a negative by a negative it makes a positive.

OpenStudy (anonymous):

I know

OpenStudy (onepieceftw):

So if you multiply 4 negatives together you're going to get a positive number

OpenStudy (anonymous):

@sicktwistd

OpenStudy (anonymous):

C and D. The number of minus signs is odd.

OpenStudy (onepieceftw):

For example: –3 • (–2) • (–4) • (–7) = 168 But if just one of those numbers is made positive the sum becomes negative.

OpenStudy (anonymous):

B?

OpenStudy (tigga281):

B is correct

OpenStudy (anonymous):

thank you just making sure

OpenStudy (anonymous):

can you explain?

OpenStudy (anonymous):

It was C

OpenStudy (whpalmer4):

yes, B is not correct because there are two negative signs, equivalent to \[(-1)*(-1)=1\]so B gives \[-2*(-7)*12*(4)=2*7*12*4=672\] answer C is correct, because \[4*(-9)*(-3)*(-1)=-36*((-3)(-1))=-36*3=-108\] odd number of negative signs in a multiplication means a negative result, unless the result is \(0\), as it is with D.

OpenStudy (tigga281):

Why did you take away my medal? At least I tried and took out my time to help complete the answer. No-one is perfect and will always get the right answer but what you did was wrong and unfriendly. @HalloimEllie

OpenStudy (whpalmer4):

medals are given to the most helpful answer, as judged by the person giving the medal. you provided no explanation, and your statement that B was correct was wrong. seems like pretty weak justification for awarding a medal, no? @TIGGA281

OpenStudy (anonymous):

@whpalmer4 On D I just counted the negative signs. Thank you for evaluating the numeric results.